For the intercontinental ballistic missile to have a range of 10500 km, what must be the speed v0 of the missile at the end of the accelaration phase? Take the distance of the accelaration phase, called the boost time, as
negligible compared to the travel time.
1
Expert's answer
2010-10-05T10:00:26-0400
As a slight acceleration phase compared to the time of flight, we can assume that the missile flies at a constant speed v0. All distance L ≈v0*T, H=g*T2/2=v0*v0/2a,
where a - is acceleration of the acceleration phase and a=g

⇒ v0 = ? ⇒ L=v0*v0g ⇒ v0 = √ (L/g)= √(10500000/9.8) ≈ 1km/s